The perimeter of a rectangle is 38 inches, and the area is 90 square inches. Find the dimensions of the rectangle.
Answer by Emsley(35)

Let length =x and breadth=y
Then xy=90 and 2x+2y=38 so x+y=19
x=90/y Sub in x+y=19 and get 90/y+y=19
Multiply both sides by y 90+y^2=19y y^2-19y+90=0
Factorising (y-10)(y-9)=0 so y=10 or 9
But y was the breadth so y=9 and x=10
That is length =10 and breadth=9